# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

### v2

@@ Abstract
DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline. In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image. This raises the question: can we trust and use all of these predictions? Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions. More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated. Based on these insights, we demonstrate that identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
Building on this viewpoint, we first address the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error. Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized. In response, we present Object-level Calibration Error (OCE), which assesses the calibration quality more effectively and is suitable for both ranking different models and identifying the most reliable predictions within a specific model. As a final contribution, we introduce a post hoc uncertainty quantification (UQ) framework that predicts the accuracy of the model on a per-image basis. By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, our framework assesses the reliability of the DETR model for each test image.
(a) Random boxes. Low Confidence.
(b) Accurate boxes. Calibrated.
@@
Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated. This figure illustrates how DETR can handle the remaining predictions (i.e., red and gray boxes). In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with accurate prediction for the bounding boxes, or (c) poorly calibrated confidence with accurate boxes. Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
I Introduction
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [1, 2, 3, 4, 5, 6] to identify and locate objects within images. More recently, DEtection TRansformer (DETR) [7] has revolutionized the field by utilizing a Transformer encoder-decoder architecture to offer a scalable end-to-end prediction pipeline, where the model predicts a set of bounding boxes and class probabilities. This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks. While notable progress has been made, the reliability of these predictions remains insufficiently investigated.
The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used. For example, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions. As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler. Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image. However, owing to DETR’s unique set-prediction mechanism, quantifying its image-level reliability is not straightforward and remains largely underexplored.
Moreover, there is an ongoing debate within the community about whether DETR can be considered an entirely end-to-end solution, free from any post-processing requirements in practice. In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used. Practitioners often employ heuristic approaches to determine this subset, such as by setting a user-defined threshold to retain only a small subset of high-confidence (e.g., $>0.7$) outputs, as seen in the official demo. Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence score. Nonetheless, how different subset selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remain under-explored, and their significance has yet to be fully recognized.
Our primary findings reveal that DETR predictions within the same image are interdependent, leading to significantly varying levels of reliability. Since DETR is trained using gradients derived exclusively from optimally matched predictions (i.e., the predictions most closely corresponding to each annotated object), it will output at least one well-calibrated prediction for each perceived object while having flexibility in handling the remaining unmatched predictions. Throughout the paper, we refer to those matched predictions as positive predictions and the others as negative predictions. In principle, to avoid an inaccurate match, DETR could assign those unmatched predictions low confidence with random bounding boxes (Figure 1a). Alternatively, it could generate similarly accurate bounding boxes with (nearly) equally calibrated confidence (Figure 1b). Otherwise, it may assign poorly calibrated confidence scores (Figure 1c). Our analysis indicates that a prediction generated by DETR mostly follows the third scheme.
Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and addresses the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability? (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set? and (RQ3) How can we accurately assess DETR’s image-level reliability? In addressing these questions, our specific contributions include:

### v3

@@ Abstract
DEtection TRansformer (DETR) and its variants have emerged as promising architectures for object detection, offering an end-to-end prediction pipeline. In practice, however, DETRs generate hundreds of predictions that far outnumber the actual objects present in an image. This raises a critical question: which of these predictions could be trusted? This is particularly important for safety-critical applications, such as in autonomous vehicles. Addressing this concern, we provide empirical and theoretical evidence that predictions within the same image play distinct roles, resulting in varying reliability levels. Our analysis reveals that DETRs employ an optimal “specialist strategy”: one prediction per object is trained to be well-calibrated, while the remaining predictions are trained to suppress their foreground confidence to near zero, even when maintaining accurate localization. We show that this strategy emerges as the loss-minimizing solution to the Hungarian matching algorithm, fundamentally shaping DETRs’ outputs. While selecting the well-calibrated predictions is ideal, they are unidentifiable at inference time. This means that any post-processing algorithm—used to identify trustworthy predictions—poses a risk of outputting a set of predictions with mixed calibration levels. Therefore, practical deployment necessitates a joint evaluation of both the model’s calibration quality and the effectiveness of the post-processing algorithm. However, we demonstrate that existing metrics like average precision and expected calibration error are inadequate for this task. To address this issue, we further introduce Object-level Calibration Error (OCE), which evaluates calibration by aggregating predictions per ground-truth object rather than per prediction. This object-centric design penalizes both retaining suppressed predictions and missed ground truth foreground objects, making OCE suitable for both evaluating models and identifying reliable prediction subsets. Finally, we present a post hoc uncertainty quantification (UQ) framework that predicts per-image model accuracy.
I Introduction
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize convolutional neural networks (CNNs) [1, 2, 3, 4, 5, 6] to identify and locate objects within images. More recently, DEtection TRansformer (DETR) [7] has revolutionized the field by utilizing a Transformer encoder-decoder architecture to offer a scalable end-to-end prediction pipeline, where the model predicts a set of bounding boxes and class probabilities. This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks. While notable progress has been made, the reliability of these predictions remains under-investigated.
Quantifying the reliability of these models’ predictions is critical across different scenarios. For instance, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions. Images where the model’s reliability is questionable may need to be reviewed by human labelers. In safety-critical applications, such as autonomous vehicles, knowledge of the reliability of a prediction in the perception pipeline for a given scene is crucial for the downstream decision-making. Without properly understanding this reliability, the system may, for example, overlook a pedestrian, leading to catastrophically wrong decisions. However, owing to DETR’s unique set-prediction mechanism, quantifying its reliability is not straightforward and remains largely underexplored.
In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used. Practitioners often employ heuristic approaches to select a subset of predictions, such as by setting a user-defined threshold to retain only a small number of high-confidence (e.g., $0.7$) outputs, as seen in the official demo [detr_demo]. Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence scores. Nonetheless, how different subset-selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remains under-explored, and their significance has yet to be fully understood.
To address this gap and move beyond heuristic subset selection, we first seek a fundamental characterization of DETRs’ predictions based on the training objective used. Our theoretical analysis shows that the Hungarian loss incentivizes a “specialist strategy”: having a single well-calibrated prediction (which we refer to as the primary prediction) to represent each ground-truth object, while suppressing all remaining predictions (referred to as the secondary predictions). This specialist strategy presents a fundamental challenge for practical deployment: while those primary predictions are trained to be reliable, they are unidentifiable at test time. This means that any post-processing algorithm poses a risk of outputting a set of predictions with mixed calibration levels. Consequently, evaluating model reliability requires jointly assessing the model’s calibration quality and the effectiveness of the post-processing algorithm in recovering the well-calibrated subset.
To facilitate this joint evaluation, we introduce object-level calibration error (OCE), which evaluates calibration along ground-truth objects rather than predictions (Section V). OCE penalizes both missing ground truth objects and retaining unreliable secondary predictions, enabling proper assessment of model + post-processing combinations, something that common existing metrics are fundamentally incapable of. For instance, average precision (AP) [16, 17, 18] favors retaining all predictions, while expected calibration error (ECE) [8, 19, 20] favors discarding nearly all predictions to achieve artificially low error by ignoring missed objects.
Finally, we leverage both our theoretical characterization of DETRs’ specialist strategy as well as OCE’s ability to identify primary and secondary predictions and develop a method for image-level uncertainty quantification. More specifically, we first demonstrate that the foreground confidence contrast between primary and secondary predictions is strongly correlated with image-level reliability. This motivates our novel framework for quantifying image-level reliability by measuring the confidence contrast between the “positive predictions” retained by an OCE post-processing and the remaining “negative” ones. We conduct numerical experiments across in-distribution, near out-of-distribution, and far out-of-distribution scenarios, and show the effectiveness of our approach (Section VI).
In short, our key contributions are:
  • We theoretically characterize and empirically validate the optimal strategy that DETR takes to minimize the Hungarian loss: it creates a single well-calibrated prediction per object while suppressing all others to near-zero foreground confidence while maintaining reasonable bounding box accuracy.
  • We introduce OCE, a metric suitable for both evaluating model calibration and identifying reliable post-processing by measuring calibration per ground-truth object, avoiding pitfalls of existing metrics.
  • We propose an algorithm to quantify image-level reliability by measuring the confidence contrast between positive and negative predictions identified by OCE. We demonstrate the effectiveness of our approach with numerical experiments on the COCO and Cityscapes datasets.
II Related Work
Calibration. Model calibration refers to how well a model’s predicted confidence scores align with the actual likelihood of those predictions being correct. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. For example, if among all instances to which the model assigns an 80% probability, the event occurs in about 80% of those instances, the model is well‑calibrated. Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
To evaluate the alignment, one can measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set-prediction nature of object detection models, it is unclear which ground-truth object corresponds to each prediction. To address these challenges, detection expected calibration error (D-ECE) [8] defines precision as the accuracy metric and matches each prediction to a ground-truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$). On the other hand, localisation-aware expected calibration error (LaECE) [19] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
Yet, when model calibration is evaluated across the entire set of predictions, which includes the secondary predictions, the resulting evaluation becomes highly biased and therefore unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in later sections, applying a threshold of $0.3$ selects the well-calibrated predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable. However, the optimal threshold is not guaranteed to be always $0.3$ for other models, like UP-DETR, which introduces potential risks. In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Uncertainty Quantification. Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, [22] proposes a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. [23] generates outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, [19] employs an auxiliary detection model capable of expressing its confidence. Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing UQ techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models. In contrast, our paper emphasizes the significance of identifying a reliable subset within the entire set of predictions for uncertainty quantification, particularly in DETRs. Further, another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
III Preliminaries
Figure 1: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are fed to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
III-A Detection Transformer (DETR)
Consider a test image $x$ and denote the set of ground-truth objects present in the image by $\mathcal{Y}$. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{Y}}_{\theta}(x)$. Each prediction $\hat{y}_{i}\in\hat{\mathcal{Y}}_{\theta}(x)$ is characterized by a bounding box $\hat{b}_{i}$ and an associated class probability distribution $\hat{p}_{i}$ over $\mathcal{C}\cup\{\emptyset\}$, where $\mathcal{C}$ is the set of foreground classes and $\emptyset$ represents the background class.
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. See Figure 1 for an illustration.
III-B Hungarian Loss in DETR
Let $\hat{\mathcal{Y}}_{\theta}(x)=\{\hat{y}_{1},...,\hat{y}_{M}\}$, where $\hat{y}_{i}=(\hat{p}_{i},\hat{b}_{i})$. This set is compared against $\mathcal{Y}=\{y_{1},...,y_{N}\}$, where $N$ is the number of ground-truth objects. To evaluate the loss, a one-to-one correspondence between the two sets is first established. The ground-truth set is padded with MNM-N background objects, and an optimal matching permutation $\hat{\sigma}$ is found by minimizing a pairwise matching cost $\mathcal{L}_{\text{match}}$ using the Hungarian algorithm [32, 7].
[…]
$\mathcal{L}(\hat{y}_{\hat{\sigma}(j)},y_{j})=\sum_{j=1}^{M}w_{j}\left[-\log\hat{p}_{\hat{\sigma}(j)}(c_{j})+\mathbbm{1}_{\{c_{j}\neq\emptyset\}}\mathcal{L}_{\text{box}}(b_{j},\hat{b}_{\hat{\sigma}(j)})\right],$ (1)
where $\hat{p}_{i}(c)$ is the predicted probability for class $c$ of the $i$-th prediction. $\mathcal{L}_{\text{box}}$ is a linear combination of the IoU and L1 loss between the two bounding boxes. This process naturally partitions the $M$ predictions into two groups for each image: the optimal positive predictions, which are matched to ground-truth foreground objects, and optimal negative predictions, which are matched to background objects. In practice, the log-probability term for negative predictions (where $c_{j}=\emptyset$) is down-weighted (e.g., by a factor of $w_{\emptyset}=0.1$):
$w_{j}=\begin{cases}w_{\emptyset}&\text{if }c_{j}=\emptyset\\ 1&\text{if }c_{j}\neq\emptyset.\end{cases}$
This structure, which penalizes high-confidence foreground predictions from negative predictions, forms the basis of our subsequent analysis.
(a) Accurate boxes. Calibrated.
(b) Random boxes. Low Confidence.
@@ III-B Hungarian Loss in DETR
Figure 2: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated. This figure illustrates how DETR handles the remaining predictions (i.e., red and gray boxes). DETR could (a) assign equally high confidence scores with accurate bounding boxes to all predictions; however, this strategy fails to optimize the Hungarian loss, as high-confidence unmatched predictions incur substantial penalties. Instead, DETR could assign low confidence to unmatched predictions, either with (b) random bounding boxes or (c) accurate bounding boxes. Our analysis reveals that DETR follows the third strategy, which explains the varying levels of reliability observed across predictions.
IV Optimal Prediction Strategies in DETR
In this work, we evaluate a model’s reliability using the concept of calibration. We define a prediction as perfectly calibrated if its predicted probability distribution matches the true posterior probability distribution of the object it is matched to. We formalize this by considering an object whose true, and potentially ambiguous, class identity is represented by a ground-truth probability distribution $p_{\text{gt}}$. A prediction is perfectly calibrated with respect to this object if its predicted distribution $\hat{p}$ exactly matches the true distribution $p_{\text{gt}}$, i.e.,
$\hat{p}(c)=p_{\text{gt}}(c)\quad\forall c\in\mathcal{C}\cup\{\emptyset\}.$ (2)
IV-A Motivation and Scope
During the inference stage at test-time, ground-truth annotations are unavailable, meaning the optimal positive predictions remain unknown. This raises the question of Which of the predictions can be trusted and used? If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 2a), the large number of predictions would not be a concern. We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
However, empirical evidence suggests that such independence is unlikely to happen, as it appears to conflict with the minimization of the Hungarian loss in Equation (1). In practice, we observe that DETR tends to assign a well-calibrated confidence score to only a single prediction per object, while suppressing remaining predictions with low foreground confidence. Intriguingly, although the Hungarian loss imposes no penalty on the localization of these background-matched optimal negative predictions—implying they could potentially output random bounding boxes (Figure 2b)—we find that DETR exhibits a strong preference for accurate localization (Figure 2c), consistently producing precise boxes even for optimal negative predictions. In the following sections, we investigate this behavior in depth.
Figure 3: Visualizations of the predictions generated by Cal-DETR. The optimal positive prediction (indexed by 0 and bordered in blue) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented. For each prediction and layer, the maximum confidence score and its corresponding label are visualized. The model’s attention mechanism enables it to learn a “specialist strategy.” When an object is confidently identified, a single primary query (the intended optimal positive prediction) is assigned a (high) calibrated confidence score. In parallel, all other remaining queries (the intended optimal negative predictions) have their confidence scores suppressed to near-zero, even while they maintain accurate bounding box predictions. This specialist strategy—where the model designates one query as the sole reliable predictor and the rest as background specialists—is demonstrably favored by the loss function over having multiple competing, moderately-confident predictions for the same object.
IV-B Exploring the Anatomy of DETR’s Predictions
We begin our qualitative analysis by visualizing and examining the outputs generated from the DETR decoders. Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward. We address this by reapplying the final feedforward network that operates on the last layer, $f_{\phi}$, to the intermediate layers. This allows us to transform each representation vector into its associated bounding box and class label. This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [34]. Sample visualizations are in Figures 3 and 4.
In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions. In this early stage, the distinction between optimal positive and negative queries can be ambiguous (e.g., Figure 3a and Figure 4a). However, the attentions through the subsequent decoder layers progressively refine these predictions. By the final layer, the model selects a single prediction and assigns a confidence score based on its understanding of the image and the object. In contrast, the confidence scores for remaining queries do not increase to the same extent as the positives and even decrease. On the other hand, in images of low reliability (where the model is uncertain), the confidence score of the optimal positive prediction does not significantly increase, while the scores of the optimal negative ones are either slightly raised or unchanged. Based on this observation, we present our main claim:
DETR learns to assign multiple predictions for each object within a given image and to designate a single prediction that is calibrated, while the remaining predictions are suppressed to have low foreground confidence, even when maintaining accurate localization. When the model’s uncertainty about its prediction (and thus about the matching process) is high, the confidence separation between those predictions diminishes.
(a) High-reliability Image
(b) Low-reliability Image
Figure 4: When the model is confident, it assigns a high confidence score to the optimal positive prediction by either increases or remains high across the decoder layers via cross-attention mechanism—thus be calibrated. While those of remaining optimal negative queries’ confidence scores are pushed to be near-zero while having accurate bounding box predictions. On the other hand, when the model is uncertain, DETR assigns a lower confidence score to the optimal positive prediction, thereby maintaining good calibration. However, conversely, it slightly increases or maintains the confidence scores for remaining predictions.
IV-C Theoretical Analysis of the Optimal Prediction Strategy
To provide theoretical support for the claim, we analyze the optimal prediction strategy in a stochastic setting. We define a random variable $y_{\text{gt}}=(c_{\text{gt}},b_{\text{gt}})$ representing the ground-truth object with distribution $Y$, where $c_{\text{gt}}\in\mathcal{C}$ represents the class that is sampled from the class distribution $p_{\text{gt}}$, and $b_{\text{gt}}$ represents the bounding box sampled from some distribution $\mathcal{B}$.
The padded set of $M$ targets is denoted by $\{y_{j}\}_{j=1}^{M}$, where $y_{1}=y_{\text{gt}}$ (the sampled object) and $y_{j}=\emptyset$ for $j\neq 1$. The model provides a set of predictions $\{\hat{y}_{i}\}_{i=1}^{M}$. The objective is to find predictions that minimize the expected total loss over the distribution $Y$:
$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\sum_{j=1}^{M}\mathcal{L}(\hat{y}_{\hat{\sigma}(j)},y_{j})\right].$
To formalize the roles of the predictions, we define the index of the optimal positive query as $p=\hat{\sigma}(1)$, while the remaining M1M-1 indices, $\{n\}_{n\neq p}$, are the optimal negative queries. Given this definition, we can decompose the total loss into terms for the optimal positive and optimal negative indices:
$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})\right]+\sum_{n\neq p}\mathcal{L}(\hat{y}_{n},\emptyset).$
The loss for the optimal positive prediction $\hat{y}_{p}=(\hat{p}_{p},\hat{b}_{p})$ matched to the $y_{\text{gt}}$ is:
$\mathbb{E}\left[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})\right]$ $=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})-\log(\hat{p}_{p}(c_{\text{gt}}))\right]$
$=\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}\left[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})\right]-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{p}(c)),$ (3)
where $\hat{p}_{p}(c)$ is the predicted probability for class $c$, as defined before. The loss for optimal negative prediction $\hat{y}_{n}$ matched to $\emptyset$ (one of the optimal negatives) is the down-weighted negative log-probability of the background class:
$\mathcal{L}(\hat{y}_{n},\emptyset)=-w_{\emptyset}\log(\hat{p}_{n}(\emptyset)).$ (4)
Here, $w_{\emptyset}$ is the down-weighting factor (e.g., 0.1) for the background class. The following result establishes a lower bound on the expected total loss.
Proposition 1.
The minimum achievable expected total loss is lower bounded by the entropy of the ground-truth class distribution, i.e.,
$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})],$
where $\mathcal{H}(p_{\text{gt}})=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$.
Proof.
See Section B-A. ∎
Next, we characterize the optimal prediction set that uniquely achieves this bound.
Proposition 2 (Optimal Prediction Set).
The unique globally optimal prediction set that minimizes $\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]$ is $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$, where